OpenStudy (anonymous):

Medall!!! The area of the square is 119 cm^2. What is the length of a side of the square?

OpenStudy (campbell_st):

what is the formula for area of a square..?

OpenStudy (anonymous):

one side squared

OpenStudy (campbell_st):

so \[Area = l^2\] where l is the side length...is that what you are saying..?

OpenStudy (anonymous):

wait sorry it would be cubed so like a = l^4 so to get l you would divide a by 4 right

OpenStudy (campbell_st):

no the formula I posted is correct... so you know Area = 119 cm^2 substitute this value \[119 = l^2\] now solve for l.... hope it makes sense

OpenStudy (anonymous):

yeah thanks!